OpenStudy (anonymous):

evaluate ln 1/sqrt e^5

OpenStudy (campbell_st):

well start by rewriting it in index form \[\frac{1}{\sqrt{e^5}} = {e^\frac{-5}{2}}\] so you are looking at \[\ln(e^{-\frac{5}{2}})\] now you can simply use log laws for the base and exponent. hope it helps

OpenStudy (anonymous):

I don't know what index form is? I don't think we've gotten that far...

OpenStudy (campbell_st):

well indexs and powers... are the same thing... but if you haven't gotten that far... you can input the question straight into a calculator to get an answer

OpenStudy (anonymous):

Awesome. Thank you.